# Generalized Kummer surfaces over finite fields

### v3

@@ 00 $1^{6},2^{4},3^{8},6^{4}$ $C_{2}$ $p\not\equiv 1\bmod 12$ $t^{4}-qt^{2}+q^{2}$
[Ar] Suppose $p\neq 2$. If a supersingular $K3$ surface $X$ is defined over ${\mathbb{F}}_{q}$ and Frobenius acts trivially on $\operatorname{NS}({\overline{X}})$, then $p^{2}$ divides $q$.
@@
Let $k={\mathbb{F}}_{p^{2r+1}}$. Assume that the order of $G$ is greater than $2$ and is not divisible by $p$. There exists a supersingular abelian surface with the Weil polynomial $f$ and a rigid action of a group $G$ if and only if $p$, $f$, and $G$ satisfy one of the following conditions.
@@
Suppose that $G$ acts on a supersingular surface $A$ over $k$. The surface $A$ is either simple or isogenous to a product of two elliptic curves, say $E_{1}$ and $E_{2}$. If $E_{1}$ and $E_{2}$ are not isogenous, and the action of $G$ is rigid, there are homomorphisms
@@
where $n$ is the order of $G$; therefore, either $G$ is trivial, or $p$ divides $n$. It follows that, $E_{1}$ and $E_{2}$ are isogenous, and
@@
According to Theorem 2.8, if $A$ is simple, then $\operatorname{End}^{\circ}(A)$ belongs to the following list:
@@
It is straightforward to check that in the last two cases, $\operatorname{End}^{\circ}(A)$ does not contain a cyclotomic field. We proved that the Weil polynomial $f=f_{A}$ of $A$ is one of the polynomials from cases $(1)-(6)$.
Let $A$ be an abelian variety with the Weil polynomial $f$. In the first two cases $L=\operatorname{End}^{\circ}(A)$ is a field:
@@
Clearly, $G$ is cyclic of order $4$ in case $(1)$, and of order $3$ or $6$ in case $(2)$. According to Corollary 3.9, in these cases there exists an abelian surface with a rigid action of $G$. By Lemmata 6.6, and 6.5, there exists an isogeny to an abelian surface with a rigid and symplectic action of $G$.
@@
If $G$ contains a cyclic subgroup of order $n\in\{5,8,12\}$, then there exists a homomorphism from ${\mathbb{Q}}(\zeta_{n})$ to ${\mathbb{H}}$. Dimension counting shows that the image contains ${\mathbb{Q}}(\sqrt{p})$, therefore $p$ divides $n$. A contradiction. According to Theorem 4.3, the group $G$ is either cyclic of order $3,4,6$, or isomorphic to $Q_{8}$, $Q_{12}$, or $\operatorname{SL}_{2}({\mathbb{F}}_{3})$.
We are going to construct rigid actions for these groups. According to Theorem 2.2, there are homomorphisms from ${\mathbb{Q}}[\zeta_{3}]$ and ${\mathbb{Q}}[\zeta_{4}]$ to ${\mathbb{H}}$. By Lemma 6.5, there exists an isogeny from $A$ to an abelian surface with a rigid and symplectic action of $G$. According to Corollary 3.9, there is an action of $Q_{8}$ or $\operatorname{SL}_{2}({\mathbb{F}}_{3})$ on some variety in the isogeny class of $A$ if and only if there is a homomorphism of ${\mathbb{H}}_{2}$ to ${\mathbb{H}}$. By Corollary 2.4, there is such a homomorphism if and only if $2$ does not split in ${\mathbb{Q}}(\sqrt{p})$. In the same way, one proves that $Q_{12}$ acts on some variety in the isogeny class of $A$ if and only if $3$ does not split in ${\mathbb{Q}}(\sqrt{p})$. According to Lemmata 6.7, and 6.8 these actions are symplectic.
@@
In particular, the argument from the proof of case $(3)$ gives the same list of groups. According to Corollary 3.9 and Theorem 2.2, there are homomorphisms from ${\mathbb{Q}}[\zeta_{3}]$ and ${\mathbb{Q}}[\zeta_{4}]$ to ${\mathbb{H}}$. We use Corollary 3.9 and Corollary 2.4 again and find that there is an action of $Q_{8}$ or $\operatorname{SL}_{2}({\mathbb{F}}_{3})$ on some variety in the isogeny class of $A$ if and only if $2$ does not split in ${\mathbb{Q}}(\sqrt{-p})$, and $Q_{12}$ acts on some variety in the isogeny class of $A$ if and only if $3$ does not split in ${\mathbb{Q}}(\sqrt{-p})$.
Cases $(5)$ and $(6)$ can be treated in the same way. ∎
@@
Assume that $k={\mathbb{F}}_{q}$, where $q=p^{2r+1}$, and $p>2$. There exists a Kummer surface $X$ over $k$ corresponding to an abelian surface $A$ with the following parameters.
@@ 00 $1^{6},2^{4},3^{8},6^{4}$ $C_{2}$ $p>2$ $t^{4}-qt^{2}+q^{2}$
If $G=C_{2}$, we apply [Ry14, Theorem 7.1]. If $p\equiv 1\bmod 4$, then, according to Proposition 7.3, there exists a supersingular surface $X$ with $\operatorname{Tr}_{X}=18$, and a supersingular surface with the trace equal to $8$.
Assume that $G=Q_{8}$, and $f_{A}(t)=(t^{2}-q)^{2}$. According to Theorem 7.10, there exists an abelian surface $A$ with the Weil polynomial $f$ and with a rigid and symplectic action of $Q_{8}$. By Lemma 7.2, the characteristic polynomial of Frobenius action on $H^{2}(\bar{A},{\mathbb{Q}}_{2})^{Q_{8}}$ is equal to $(t-q)^{2}(t+q)$. We construct an abelian surface $B$ with a $2$-isogeny $B\to A$ and an action of $Q_{8}$ such that each orbit of the action of Frobenius on $B[2]({\bar{k}})$ is a subset of an orbit of $Q_{8}$. According to Proposition 5.5 and Lemma 7.1, the zeta function of the Kummer surface $K(B,Q_{8})$ is $(1^{21},2)$.
The algebra $\operatorname{End}^{\circ}(A)$ is generated by ${\mathbb{H}}_{2}\cong{\mathbb{Q}}[Q_{8}]^{\mathrm{rig}}$ over its center ${\mathbb{Q}}(\sqrt{p})$. Let $i\in{\mathbb{H}}_{2}$ be the image of an element of $Q_{8}$ of order $4$, then there is a relation $(iF)^{2}=-q$. Let $T$ be a free submodule of $T_{2}(A)$ over the algebra $R={\mathbb{Z}}_{2}[i,y]$, where $y=(1+iF)/2$. Note that $T$ is $G$ invariant. According to Lemma 2.5, there exists $B$ such that $T_{2}(B)\cong T$. In particular, $F$ is equal to $i$ on $B[2]({\bar{k}})$. ∎

### v4

@@ 0 $1^{6},2^{4},3^{8},6^{4}$ $C_{2}$ $p\not\equiv 1\bmod 12$ $t^{4}-qt^{2}+q^{2}$
[Ar74] Suppose $p\neq 2$. If a supersingular $K3$ surface $X$ is defined over ${\mathbb{F}}_{q}$ and Frobenius acts trivially on $\operatorname{NS}({\overline{X}})$, then $p^{2}$ divides $q$.
@@
Let $k={\mathbb{F}}_{p^{2r+1}}$. Assume that the order of $G$ is greater than 22 and is not divisible by pp. There exists a supersingular abelian surface with the Weil polynomial $f$ and a rigid action of a group $G$ if and only if $p$, $f$, and $G$ satisfy one of the following conditions.
@@
Suppose that $G$ acts on a supersingular surface AA over $k$. The surface AA is either simple or isogenous to a product of two elliptic curves, say $E_{1}$ and $E_{2}$. If $E_{1}$ and $E_{2}$ are not isogenous, and the action of $G$ is rigid, there are homomorphisms
@@
where $n$ is the order of $G$; therefore, either $G$ is trivial or $p$ divides $n$. It follows that $E_{1}$ and $E_{2}$ are isogenous, and
@@
According to Theorem 2.8, if AA is simple, then $\operatorname{End}^{\circ}(A)$ belongs to the following list:
@@
It is straightforward to check that in the last two cases, $\operatorname{End}^{\circ}(A)$ does not contain a cyclotomic field. We proved that the Weil polynomial $f=f_{A}$ of AA is one of the polynomials from cases (1)(6)(1)-(6).
Let AA be an abelian variety with the Weil polynomial $f$. In the first two cases $L=\operatorname{End}^{\circ}(A)$ is a field:
@@
Clearly, $G$ is cyclic of order 44 in case $(1)$, and of order 33 or 66 in case $(2)$. According to Theorem 1.3, in these cases there exists an abelian surface with a rigid action of $G$. By Lemmata 6.6, and 6.5, there exists an isogeny to an abelian surface with a rigid and symplectic action of $G$.
@@
If $G$ contains a cyclic subgroup of order $n\in\{5,8,12\}$, then there exists a homomorphism from ${\mathbb{Q}}(\zeta_{n})$ to $\mathbb{H}$}. Dimension counting shows that the image contains ${\mathbb{Q}}(\sqrt{p})$, therefore $p$ divides $n$. A contradiction. According to Theorem 4.3, the group $G$ is either cyclic of order $3,4,6$ or isomorphic to $Q_{8}$, $Q_{12}$, or $\operatorname{SL}_{2}({\mathbb{F}}_{3})$.
We are going to construct rigid actions for these groups. According to Theorem 2.2, there are homomorphisms from ${\mathbb{Q}}[\zeta_{3}]$ and ${\mathbb{Q}}[\zeta_{4}]$ to $\mathbb{H}$}. By Lemma 6.5, there exists an isogeny from AA to an abelian surface with a rigid and symplectic action of $G$. According to Theorem 1.3, there is an action of $Q_{8}$ or $\operatorname{SL}_{2}({\mathbb{F}}_{3})$ on some variety in the isogeny class of AA if and only if there is a homomorphism of ${\mathbb{H}}_{2}$ to $\mathbb{H}$}. By Corollary 2.4, there is such a homomorphism if and only if 22 does not split in ${\mathbb{Q}}(\sqrt{p})$. In the same way, one proves that $Q_{12}$ acts on some variety in the isogeny class of AA if and only if 33 does not split in ${\mathbb{Q}}(\sqrt{p})$. According to Lemmata 6.7, and 6.8 these actions are symplectic.
@@
In particular, the argument from the proof of case $(3)$ gives the same list of groups. According to Theorem 1.3 and Theorem 2.2, there are homomorphisms from ${\mathbb{Q}}[\zeta_{3}]$ and ${\mathbb{Q}}[\zeta_{4}]$ to $\mathbb{H}$}. We use Corollary 1.3 and Corollary 2.4 again and find that there is an action of $Q_{8}$ or $\operatorname{SL}_{2}({\mathbb{F}}_{3})$ on some variety in the isogeny class of AA if and only if 22 does not split in ${\mathbb{Q}}(\sqrt{-p})$, and $Q_{12}$ acts on some variety in the isogeny class of AA if and only if 33 does not split in ${\mathbb{Q}}(\sqrt{-p})$.
The cases $(5)$ and $(6)$ can be treated in the same way. ∎
@@
Assume that $k={\mathbb{F}}_{q}$, where $q=p^{2r+1}$, and $p>2$. There exists a Kummer surface $X$ over $k$ corresponding to an abelian surface AA with the following parameters.
@@ 0 $1^{6},2^{4},3^{8},6^{4}$ $C_{2}$ $p>2$ $t^{4}-qt^{2}+q^{2}$
If $G=C_{2}$, we apply [Ry14, Theorem 7.1]. If $p\equiv 1\bmod 4$, then, according to Proposition 7.3, there exists a supersingular surface $X$ with $\operatorname{Tr}_{X}=18$, and a supersingular surface with the trace equal to 88.
Assume that $G=Q_{8}$, and $f_{A}(t)=(t^{2}-q)^{2}$. If $p\equiv 3(4)$, then, according to Theorem 7.10, there exists an abelian surface AA with the Weil polynomial $f$ and with a rigid and symplectic action of $Q_{8}$. By Lemma 7.2, the characteristic polynomial of Frobenius action on $H^{2}(\bar{A},{\mathbb{Q}}_{2})^{Q_{8}}$ is equal to $(t-q)^{2}(t+q)$. We construct an abelian surface $B$ with a 22-isogeny $B\to A$ and an action of $Q_{8}$ such that each orbit of the action of Frobenius on $B[2]({\bar{k}})$ is a subset of an orbit of $Q_{8}$. According to Proposition 5.5 and Lemma 7.1, the zeta function of the generalized Kummer surface $K(B,Q_{8})$ is equal to $(1^{21},2)$.
The algebra $\operatorname{End}^{\circ}(A)$ is generated by ${\mathbb{H}}_{2}\cong{\mathbb{Q}}[Q_{8}]^{\mathrm{rig}}$ over its center ${\mathbb{Q}}(\sqrt{p})$. Let $i\in{\mathbb{H}}_{2}$ be the image of an element of $Q_{8}$ of order 44, then there is a relation $(iF)^{2}=-q$. Let $T$ be a free submodule of $T_{2}(A)$ over the algebra $R={\mathbb{Z}}_{2}[i,y]$, where $y=(1+iF)/2$. Note that $T$ is $G$ invariant. According to Lemma 2.5, there exists $B$ such that $T_{2}(B)\cong T$. In particular, $F$ is equal to $i$ on $B[2]({\bar{k}})$. ∎